### v2

11institutetext: Center for Scientific Computing - sciCORE, University of Basel, Klingelberstrasse 61, 4056 Basel, Switzerland
11email: ruben.cabezon@unibas.ch
22institutetext: Center for Data Analytics - CeDA, University of Basel, Klingelberstrasse 61, 4056 Basel, Switzerland 33institutetext: Departament de Física. Universitat Politècnica de Catalunya (UPC). Av. Eduard Maristany 16, 08019 Barcelona, Spain
33email: domingo.garcia@senior.upc.edu
44institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), 08860 Castelldefels (Barcelona), Spain 55institutetext: Universität Heidelberg, Zentrum für Astronomie, Institut für Theoretische Astrophysik, Albert-Ueberle-Str. 2, 69120 Heidelberg, Germany 66institutetext: Department of Mathematics and Computer Science, University of Basel, Spiegelgasse 1, 4051 Basel, Switzerland 77institutetext: Swiss National Supercomputing Center (CSCS), Via Trevano 131, 6900 Lugano, Switzerland 88institutetext: Department of Astrophysics, University of Zürich, Winterthurerstrasse 190, 8057 Zürich, Switzerland 99institutetext: Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Im Neuenheimer Feld 225, 69120 Heidelberg, Germany
From particles to precision
Simulating subsonic turbulence with Smoothed Particle Hydrodynamics
Rubén M. Cabezón    Domingo García-Senz    Oliver Avril    Osman Seckin Simsek    Sebastian Keller    Axel S. Lechuga    Lucio Mayer    Ralf Klessen    Florina M. Ciorba
(Received September 30, 20XX)
Direct numerical simulation of subsonic turbulence with smoothed particle hydrodynamics (SPH) has traditionally been hampered by zeroth-order (E0) errors, inaccurate gradient evaluations, and excessive numerical dissipation. We demonstrate that the recently developed SPH-EXA code can overcome these challenges, producing results similar to those of other state-of-the-art non-SPH codes, such as AREPO and GIZMO. We show that SPH-EXA accurately reproduces the Kolmogorov inertial range scaling in the subsonic regime with increasing resolution, closely matching state-of-the-art non-SPH methods. This is an outstanding result given the important role of SPH codes in astrophysics and cosmology, and the ubiquity of turbulent phenomena.
Key Words.:
hydrodynamics – turbulence – Methods: numerical
[…]
1 Introduction
Smoothed Particle Hydrodynamics (SPH) is a mesh-free, conservative method for continuum mechanics widely used in many different fields since its conception around 1980 (Lucy, 1977; Gingold and Monaghan, 1977). Its strengths include excellent conservation properties and a natural treatment of fully three-dimensional flows (Liu and Liu, 2010; Price, 2012b). Despite these advantages, SPH has historically struggled in some regimes, most notably in accurately simulating subsonic turbulence.
Turbulence is a fundamental multiscale phenomenon in fluid dynamics, characterized by chaotic and nonlinear motion (Frisch, 1995). It occurs in both natural environments and technological applications, making its study central across science and engineering. In astrophysics, turbulence influences processes ranging from accretion flows to the interstellar medium, shaping star formation, galaxy evolution, and the emergence of large-scale structure. Despite its ubiquity and importance, turbulence remains one of the most challenging phenomena to model and predict because of its complex, non-linear nature. Progress in this field beyond the heuristic but very successful pioneering model of Kolmogorov (Kolmogorov, 1941) can lead to advances in technology, improved environmental management, and deeper insight into fundamental processes of nature.
The direct numerical simulation (DNS) of turbulence requires high-resolution and state-of-the-art codes. In general, there are no significant discrepancies between SPH simulations and grid-based methods for supersonic turbulence (Kitsionas et al., 2009). Nevertheless, reproducing subsonic or transonic turbulence with SPH has traditionally been a challenge because it failed to produce results with sufficient quality to compete with grid-based methods for a similar resolution (Bauer and Springel, 2012). The first studies of subsonic turbulence with SPH in the context of cosmological simulations (Dolag et al., 2005; Valdarnini, 2011) found rough agreement with the Kolmogorov power spectra but only in a very limited range of wavenumbers. More recent DNS attempts led to minor (Bauer and Springel, 2012) or moderate improvements (Price, 2012a), with the latter highlighting the importance of reducing the viscosity away from shocks after showing that the Reynolds number is proportional to the Mach number in SPH simulations. However, the velocity power spectra still showed significant differences with those obtained with mesh-based calculations. All of these studies identified zeroth-order errors ($E0$), excessive dissipation, and inaccurate gradient estimation as the main causes of the difficulties in simulating turbulence with SPH in low-Mach numbers.
In this work, we present SPH simulations of subsonic turbulence with SPH-EXA [^1] https://github.com/sphexa-org/sphexa
https://deepwiki.com/sphexa-org/sphexa
(Cavelan et al., 2020; Keller et al., 2023), a GPU-native state-of-the-art and highly scalable hydrodynamic code that incorporates the latest advances in high performance Computing (HPC), such as efficient domain decomposition, GPU-based neighbor search, on-the-fly initial conditions generation, and in-situ visualization. It also incorporates the latest improvements in the SPH technique, such as an integral approach to calculate gradients, generalized volume elements, and a slope-limited reconstruction (SLR) of the velocity field to control dissipation, among other improvements (see Appendices A and B). Equipped with such a set of resources, we show that the simulations of subsonic turbulence with SPH-EXA can compete with those obtained with top reference codes such as AREPO (Springel, 2010) and GIZMO (Hopkins, 2015). Our results show that, as resolution increases, we recover a growing portion of the inertial range of the Kolmogorov cascade. Owing to the excellent scaling properties of SPH-EXA, this will allow DNS to probe an ever larger fraction of the turbulent inertial range in the coming years.
Figure 1: Comparison of the power spectra of the velocity field with different elements incorporated into the SPH solver. The dashed line is the theoretical Kolmogorov power spectrum of $\sim k^{-5/3}$. STD denotes standard SPH and is taken from (Bauer and Springel, 2012, simulation S3 in their Fig. 6, with $256^{3}$ particles). All other simulations were performed with SPH-EXA and $250^{3}$ particles. The meaning of the remaining acronyms is provided in Sect. 3.
Figure 2: Comparison of the instantaneous power spectra of the velocity field at $t=10$ at increasing resolution. Each line corresponds to an increase of a factor of 2 in resolution with respect to the previous curve.
2 Initial conditions
We have designed an embedded initial conditions (IC) generator in SPH-EXA. In all our tests, we use a template of a relaxed box with $50^{3}$ particles. This box has previously been relaxed to achieve a glass-like configuration (Arth et al., 2019) with constant density and periodic boundary conditions. Replicating this box in different directions allows us to create on-the-fly initial conditions with a user-defined number of particles and in a distributed manner, only limited by the memory of the system.
All simulations use a periodic box of side $L=1$ with uniform initial density $\rho=1$ and pressure $P=1$, and ’quasi’-isothermal equation of state ($\gamma=1.001$). Turbulence is driven by stochastic forcing modes at different wavelengths k in the range $1<|\mathbf{k}|<3$, with amplitudes following a parabolic distribution $\propto 1-(k-k_{c})^{2}$ peaking at $k_{c}=2$, ensuring that kinetic energy is injected only at the largest scales of the spectrum. The resulting forcing is built from a mixture of $1/3$ compressive and $2/3$ solenoidal modes, whose phases evolve following a stochastic Ornstein-Uhlenbeck process (Uhlenbeck and Ornstein, 1930; Federrath et al., 2010). Simulations start from homogenous ICs and are driven for $t=10$ (in units of sound-crossing time, $t_{sc}=L/c=1$), until a statistically stationary state with RMS Mach number $\mathcal{M}\sim 0.34$ is reached.
3 Results
We present the main results of the simulations carried out with SPH-EXA. Technical details, such as the integral approach to derivatives (IA) used to evaluate gradients, the use of generalized volume elements (GVE), the treatment of dissipation with both variable AV coefficients (VAV), and slope reconstruction technique (SLR), are deferred to the appendices. The appendices also include additional diagnostics based on the second-order velocity structure function, highlighting the advantages of SPH’s fully Lagrangian description.
3.1 Impact of solver elements at fixed resolution
To evaluate the weight and significance of the different elements that contribute to our SPH implementation, we conducted identical simulations while progressively incorporating more of these elements. Figure 1 shows the power spectrum of the velocity field of simulations of $250^{3}$ particles with increasing complexity, compared to the S3 simulation of (Bauer and Springel, 2012) with $256^{3}$ particles and the classic SPH formulation (i.e. gradients calculated directly as kernel derivatives, no GVEs, and no viscosity switches), labeled STD in our plot.
Figure 3: Slice of thickness $2h$ centered on $z=0$ showing the velocity magnitude with increasing resolution (SPH-EXA). We represent particles directly.
The inclusion of IA recovers the correct slope in the inertial range of the Kolmogorov cascade at low $k$, considerably extending it compared to STD (Valdarnini, 2016), but it does not remove the overly steep slope $\sim-4$ at large $k$. At these small scales something else is needed. Adding GVE and especially VAV moves in the right direction, revealing the bottleneck effect as an energy excess near the resolution limit, as commonly seen in mesh codes. Finally, the slope-limited reconstruction (SLR) provides the decisive improvement: it extends the inertial range almost a decade relative to STD, makes the bottleneck clearly visible, and yields a dissipative-range slope comparable to non-Lagrangian codes (see Fig. 5).
3.2 Resolution scaling and flow morphology
In Fig. 2 we present the power spectrum of the velocity field for simulations performed with SPH-EXA and with different particle counts at the end of each simulation ($t=10$). It is clear that with increasing resolution, the inertial range over the Kolmogorov scale is extended to larger wavenumber, and more importantly, that the rate of enhancement with increasing number of particles is not marginal or small, but considerable and consistent with that obtained with modern mesh-based and moving-mesh methods. This is a novel result, which means that SPH can indeed resolve smaller scales with higher resolution, in opposition to the results found with older SPH implementations (e.g., Bauer and Springel, 2012). Figure 3 shows a thin slice of the velocity field for increasing resolution with SPH-EXA, from $200^{3}$ up to $3200^{3}$, where it is evident how we can resolve a delicate and rich mixture of large- and small-scale features. Figure 4 compares the mass-weighted and volume-weighted density PDFs, with a fitted normal distribution. We can see the expected tail bias towards low densities, since in subsonic flows, local pressure increases in colliding flows can prevent the buildup of high-density regions (Konstandin et al., 2012).
Figure 4: Instantaneous density PDF for the $400^{3}$ simulation at $t=10$. Comparison of mass-weighted and volume-weighted PDFs.
3.3 Cross-code comparison
We can now compare our results directly to other state-of-the-art codes: moving-mesh and meshless finite-volume methods. Figure 5 shows velocity power spectra from AREPO, GIZMO, and SPH-EXA. All three reproduce a comparable inertial range and a bottleneck of similar magnitude at comparable $k$. This suggests that, once the injection method and effective viscosity are matched, the cascade between the forcing and the dissipation wavenumbers behaves similarly across methods. Note that we compare runs with a similar number of fluid elements, which does not necessarily imply the same spatial resolution. In SPH, resolution is set by the smoothing length, which covers significantly more than just one particle. At $256^{3}$ cells, a proxy is the cell size, which for AREPO is $\sim 1/256\sim 3.9\times 10^{-3}$, comparable to our $400^{3}$ SPH simulation, with an average of $h\sim 3.6\times 10^{-3}$. Spatial resolution matters the most in the dissipation regime, where AREPO, GIZMO, and SPH-EXA have different transfer functions. In particular, GIZMO and SPH-EXA show a similar gentler slope in the roll-off, while AREPO is slightly steeper, pointing to a stronger high-$k$ damping. However, these differences are modest and the qualitative behavior persists at higher resolution, with the break shifting to larger $k$, as expected.
Figure 5: Comparison of the velocity power spectra for different codes at $t=10$. AREPO and GIZMO curves are from their respective papers.
Figure 6 shows, from top to bottom, the module of the velocity, density, and enstrophy of the fluid at the end of our simulations. The two columns on the left are an excerpt of the results with AREPO and the SPH used in (Bauer and Springel, 2012). It is clear from these two columns that the SPH used in that work was unable to resolve subsonic turbulence because it lacks all small-scale features. The column on the right shows the results of SPH-EXA with $250^{3}$ particles. The similarities with the results of AREPO are evident, and there is a rich mix of fine structures. This can also be seen in Fig. 3, which illustrates a slice of the module of the velocity field at $t=10$ for different resolutions. SPH-EXA also produces an excellent second-order velocity structure function (appendix C). Additional colormaps of the highest resolution simulation, $3200^{3}$, can be found in appendix D.
Figure 6: Comparison of SPH-EXA results with AREPO and SPH-tav in Bauer and Springel (2012) at $t=25$. From top to bottom: module of velocity, density, and enstrophy. The two columns on the left are an excerpt of the original results of AREPO and their SPH with reduced artificial viscosity (both with $256^{3}$ fluid elements) of Bauer and Springel (2012) shown in their Fig. 4. The column on the right is the result of SPH-EXA with $250^{3}$ particles. The fields show directly particles in a thin slice around $Z=0$. No rasterization was performed.
4 Summary and conclusions
SPH-EXA combines a modern SPH formulation with state-of-the-art parallelization and GPU acceleration. This enables simulations that mitigate historical SPH limitations, such as excessive dissipation and inaccurate gradient evaluations, that have long hindered SPH from matching the performance of grid-based and moving-mesh codes in this field. Our findings not only align well with results from non-SPH codes like AREPO and GIZMO, but also illustrate the robustness of SPH in handling complex turbulent flows at resolutions previously thought unattainable with this method.
We demonstrate that, with increasing resolution, the inertial range of the power spectra extends to larger wavenumbers, exhibiting features comparable to those seen in non-SPH codes. These improvements allow SPH to better exploit its intrinsic strengths: truly Lagrangian sampling (particles are the fluid elements, avoiding tracer advection and particle-to-grid interpolation noise), excellent conservation properties (exact mass advection and pairwise-conservative exchanges of momentum and energy), and flexibility across physical regimes. In particular, SPH couples naturally and accurately to tree methods for self-gravity, supporting its use in more demanding astrophysical contexts.
Acknowledgements.
This work has been supported by the Swiss Platform for Advanced Scientific Computing (PASC) project ”SPH-EXA: Optimizing Smoothed Particle Hydrodynamics for Exascale Computing”. It has also been carried out as part of the SKACH consortium through funding from SERI. We also acknowledge the financial support of the European Research Council through the ERC Synergy Grant “ECOGAL” (project ID 855130), from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) “STRUCTURES”, and from the German Ministry for Economic Affairs and Climate Action in project “MAINN” (funding ID 50OO2206). The numerical calculations have been supported by the EuroHPC-JU Extreme Scale Access Mode through the project ”TGSF: The role of turbulence and gravity in star formation” (EHPC-EXT-2023E01-031). The authors acknowledge the support of the LUMI Supercomputer Center, the Swiss National Supercomputing Center (CSCS - allocations c32 and ch18), and the Center for Scientific Computing (sciCORE) at the University of Basel.
Appendix A SPH Ingredients
In this section, we discuss the ingredients of our SPH formulation that have the largest impact on our simulations of subsonic turbulence. Namely, the calculation of gradients, dissipation, and generalized volume elements.
A.1 Gradients
SPH-EXA implements an integral approach to calculate derivatives (García-Senz et al. 2012; Cabezón et al. 2012).
The IA formalism calculates gradients of an arbitrary function $f$ by solving the following matrix equation:
$\left[\begin{array}[]{c}\partial f/\partial x_{1}\\ \partial f/\partial x_{2}\\ \partial f/\partial x_{3}\\ \end{array}\right]_{a}=\left[\begin{array}[]{ccc}\tau_{11}&\tau_{12}&\tau_{13}\\ \tau_{21}&\tau_{22}&\tau_{23}\\ \tau_{31}&\tau_{32}&\tau_{33}\end{array}\right]^{-1}_{a}\left[\begin{array}[]{c}I_{1}\\ I_{2}\\ I_{3}\\ \end{array}\right]_{a}\,$, (1)
where the $\tau$ tensor elements are defined as the weighted sum of the dyadic product of the relative distances $(\mathbf{x}_{b}-\mathbf{x}_{a})\otimes(\mathbf{x}_{b}-\mathbf{x}_{a})$:
$\tau_{ij,a}=\sum_{b}V_{b}\penalty 10000\ (x_{i,b}-x_{i,a})(x_{j,b}-x_{j,a})\penalty 10000\ W_{ab}(h_{a});\penalty 10000\ i,j=1,3\,.$ (2)
Here $x_{i}$ is the vector position coordinates, $V_{b}$ the volume element, $W_{ab}$ the interpolation kernel, and $I_{i,a}$ the vector integral elements, which are calculated as:
$I_{i,a}=\left[\sum_{b}V_{b}f_{b}\penalty 10000\ (x_{i,b}-x_{i,a})\penalty 10000\ W_{ab}(h_{a})\right];\penalty 10000\ i=1,3\,.$ (3)
As a consequence, the gradient calculation is:
$\nabla_{i}f_{a}=\sum_{b}V_{b}(f_{b}-f_{a})A_{i,ab}(h_{a});\penalty 10000\ i=1,3\,$, (4)
where,
$A_{i,ab}(h_{a})=\sum_{j=1}^{3}c_{ij,a}(h_{a})(x_{j,b}-x_{j,a})W_{ab}(h_{a})\,$, (5)
where the summation is up to 3 for 3D simulations and $c_{ij,a}$ the elements of the inverse of tensor $\mathcal{T}$, that is, $\mathcal{C}=\mathcal{T}^{-1}$ (Eq. 1).
A.2 Dissipation
The basic dissipation control in SPH-EXA is done using the following artificial viscosity (AV) prescription:
$\Pi_{ab}^{AV}=\begin{cases}-\frac{\alpha}{2}v_{ab}^{sig}\penalty 10000\ w_{ab}&\text{for\penalty 10000\ \penalty 10000\ $\mathbf{x}_{ab}\cdot\mathbf{v}_{ab}<0$}\\ 0&\text{otherwise}\end{cases}\,$, (6)
where $\alpha$ is the viscous parameter and is set to 1. The signal velocity estimator between a pair of particles is:
$v^{sig}_{ab}=\frac{1}{2}(c_{a}+c_{b})-2w_{ab}\,$, (7)
Finally, $c_{a}$ and $c_{b}$ are the local speed of sound for each particle, and $w_{ab}=(\mathbf{v}_{ab}\cdot\mathbf{x}_{ab})/|\mathbf{x}_{ab}|$.
A linear velocity field implies a smooth flow without the presence of shocks. Therefore, linear velocity fields should have a vanishing AV. We reconstruct the local velocity field for each particle by removing the linear component of the velocity jump that is used to trigger and calculate the AV in Eq. 6 (Frontiere et al. 2017). This is the slope-limited reconstruction (SLR) method. Assuming that it is differentiable in the neighborhood of each SPH particle, the best linear approximation of the velocity field is given by multiplying its Jacobian by the distance to its neighbors. Instead of using $\mathbf{v}_{ab}=\mathbf{v}_{a}-\mathbf{v}_{b}$ to calculating $w_{ab}$, we use $\mathbf{v}^{\prime}_{ab}=\mathbf{v}^{\prime}_{a}-\mathbf{v}^{\prime}_{b}$, defined as:
$v^{\prime}_{a,i}$ $\equiv v_{a,i}-\frac{1}{2}\phi_{ab}\mathbf{J}_{\mathbf{v}_{a}}\mathbf{x}^{T}_{ab}$ (8)
$v^{\prime}_{b,i}$ $\equiv v_{b,i}+\frac{1}{2}\phi_{ba}\mathbf{J}_{\mathbf{v}_{b}}\mathbf{x}^{T}_{ab}$ (9)
where $\mathbf{x}^{T}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b}$ is the column vector of relative distance and $\mathbf{J}_{\mathbf{v}}$ is the Jacobian of the velocity field. In the case of discontinuities, such a linear estimator is not accurate enough; hence, the $\phi_{ab}$ factor is designed to go to zero in these cases, while it is kept close to 1 when the field is smooth. To this end, we use a van Leer-like limiter (Frontiere et al. 2017):
$\phi_{ab}$ $=\max\left[0,\min\left(1,\frac{4F_{ab}}{(1+F_{ab})^{2}}\right)\right]\times\kappa_{ab}\,$, (10)
[…]
$q_{ab}$ $\equiv\min(q_{a},q_{b})\,$, (12)
$q_{crit}$ $\equiv\left(\frac{32\pi}{3nb_{a}}\right)^{1/3}\,$, (13)
$F_{ab}$ $\equiv\frac{\mathbf{x}_{ab}\mathbf{J}_{\mathbf{v}_{a}}\mathbf{x}^{T}_{ab}}{\mathbf{x}_{ab}\mathbf{J}_{\mathbf{v}_{b}}\mathbf{x}^{T}_{ab}}\,$, (14)
where $q_{a}$ is the normalized distance between neighbors ($q_{a}=|\mathbf{x}_{b}-\mathbf{x}_{a}|/h_{a}$), $q_{crit}$ is the average of the interparticle distance normalized to the smoothing length in 3D, and $nb_{a}$ is the number of neighbors of particle $a$. The parameter $q_{fold}$ determines the width of the correction factor $\kappa_{ab}$ (Eq. 11) that is applied to the van Leer limiter (Eq. 10) and we use the recommended value (Frontiere et al. 2017) of $q_{fold}=0.2$.
The final step is to control excessive dissipation effectively. This can be done by allowing a dynamic individual viscous parameter $\alpha_{a}$ in Eq. 6, rather than having a constant value for all SPH particles. Several methods have been developed to calculate $\alpha_{a}$, but the nominal works of Cullen and Dehnen (2010) and Read and Hayfield (2012) are the inspiration behind our implementation in SPH-EXA. The main idea is that $\alpha_{a}$ should be initialized by default at a rather low value (usually $\alpha_{a}=0.05$). Then it increases rapidly if dissipation is needed and decays exponentially to its original value when it is not needed, effectively acting as a switch for the AV at the position of each SPH particle.
We start by calculating a local viscous parameter $\alpha_{loc,a}$, following the definition (Read and Hayfield 2012):
[…]
where $\mathcal{D}_{a}=h^{2}_{a}|\nabla(\nabla\cdot\mathbf{v}_{a})|$ predicts flow convergence by comparing the gradient of $\nabla\cdot\mathbf{v}_{a}$ with the local value of $\nabla\cdot\mathbf{v}_{a}$. If the former is large enough and the fluid is actually converging ($\nabla\cdot\mathbf{v}_{a}<0$), $\alpha_{loc,a}$ adopts a value between 0 and $\alpha_{max}$. Furthermore, the term $0.05c_{a}$, where $c_{a}$ stands for the local speed of sound, acts as a base value for the size of the velocity fluctuations that can trigger dissipation, and it also works as a term to prevent divergent $\alpha_{loc,a}$.
Once we know $\alpha_{loc,a}$ we have two options:
  • if $\alpha_{loc,a}\geq\alpha_{a}$: we instantaneously rise the dissipation setting $\alpha_{a}=\alpha_{loc,a}$ ,
  • if $\alpha_{loc,a}<\alpha_{a}$: we let $\alpha_{a}$ decay smoothly .
In the case that $\alpha_{a}$ has to decay, we evaluate an instantaneous decay rate $\dot{\alpha}_{a}$ as in (Read and Hayfield 2012):
$\dot{\alpha}_{a}=\begin{dcases}(\alpha_{loc,a}-\alpha_{a})/\tau_{a}&\text{if\penalty 10000\ \penalty 10000\ $\alpha_{min}<\alpha_{loc,a}<\alpha_{a}$}\,,\\ (\alpha_{min}-\alpha_{a})/\tau_{a}&\text{if\penalty 10000\ \penalty 10000\ $\alpha_{min}\geq\alpha_{loc,a}$}\,.\end{dcases}$ (16)
The decay time $\tau_{a}$ is defined as $\tau_{a}=h_{a}/v_{sig,max}$, and it is a function of the local spatial resolution and the maximum signal velocity $v_{sig,max}$ in the neighborhood of particle $a$, calculated according to Eq. 7. The effect of Eq. 16 is to allow alpha to decay exponentially to $\alpha_{loc,a}$, or $\alpha_{min}$ if $\alpha_{loc,a}$ is too low. In this way, we ensure a minimum dissipation that helps reduce the random noise that might arise from particle disorder. However, we note that even though all particles have at least a minimum value $\alpha_{min}$, dissipation is automatically cut off when $\mathbf{r}_{ab}\cdot\mathbf{v}_{ab}\geq 0$, as stated in Eq. 6. Also, there is no restriction for $\alpha_{min}$, which can be set to zero.
In this respect, it is worth noticing the latest works pointing to the interesting possibility of not needing switches at all, if a modulated version of the SLR is used (Price 2024; Sandnes et al. 2025; García-Senz and Cabezón 2025). This is achieved by taking constant $\alpha=1$ and multiplying the limiter $\phi_{ab}$ in Eqs. 8 and 9 by:
$1-\mathcal{B}^{p}\,$, (17)
where $\mathcal{B}$ is the Balsara limiter,
${\mathcal{B}}_{a}=\frac{|\nabla\cdot\mathbf{v}|_{a}}{|\nabla\cdot\mathbf{v}|_{a}+|\nabla\times\mathbf{v}|_{a}+10^{-4}\frac{c_{a}}{h_{a}}}\,.$ (18)
García-Senz and Cabezón (2025) showed that taking $p=2$ in Eq. 17 provides the best results across a series of tests involving shocks and/or shear, improving considerably over the use of switches and switches+SLR techniques. However, in order to get the best results in simulating subsonic turbulence, they recommend including an additional control in the amplitude of the AV to avoid spurious triggering of dissipation by the random motion of particles in the smaller scales. This can be done either with a switch, as shown before, or directly with a limiter, such as Balsara.
A.3 Volume elements
In the context of SPH, where we are discretizing the SPH integral interpolant, it becomes necessary to compute the volume element that corresponds to each particle. An uncomplicated yet highly effective method is $dx^{\prime 3}\equiv V_{b}\sim m_{b}/\rho_{b}$. However, it is known (Saitoh and Makino 2013; Hopkins 2013) that the discretization of the volume element, present in most SPH equations, can be improved beyond this classical approximation using generalized volume elements (GVE).
GVE are usually expressed as:
$V_{a}=\frac{X_{a}}{\sum_{b=1}^{nb}X_{b}W_{ab}}\,$, (19)
where $X$ is an arbitrary function. In SPH-EXA, we adopted the GVE described in the work of García-Senz et al. (2022), which has been shown to provide a better treatment of discontinuities, stronger resistance to tensile instability, better partition of unity, and more accurate gradient evaluation:
$X_{a}=\frac{m_{a}}{\rho^{0}_{a}}\,$, (20)
where $\rho^{0}_{a}=\sum_{b}m_{b}W_{ab}$ is the standard SPH density calculation. With Eqs. 19 and 20 we can simply calculate the particle density as $\rho_{a}=m_{a}/V_{a}$.
A.4 SPH-EXA hydrodynamic equations
The final set of hydrodynamic equations in SPH-EXA is as follows:
$\rho_{a}=$ $m_{a}/V_{a}\,$, (21)
[…]
$\begin{split}\frac{du_{a}}{dt}=&\frac{X_{a}^{2-\sigma}P_{a}}{m_{a}^{2}\Omega_{a}k_{a}}\sum_{b}\sum^{d}_{i=1}m_{b}X_{b}^{\sigma}\left[(\mathbf{v}_{a}-\mathbf{v}_{b})\cdot\mathbf{A}_{ab}(h_{a})\right]+\left(\frac{du_{a}}{dt}\right)^{AV}\,.\end{split}$ (23)
Here $P_{a}$ is the pressure, $X_{a}=m_{a}/\rho^{0}_{a}$ the selected function for the GVE, $k_{a}$ the normalization of the GVE ($k_{a}\equiv\sum_{b}X_{b}W_{ab}$ in Eq. 19), $A_{i,ab}(h_{a})$ the IA term, and $\Omega_{a}$ the grad-h term. The parameter $\sigma$ controls how close the implementation is to a pure Lagrangian scheme (see García-Senz et al. 2022). The last term in the energy equation includes the contributions of artificial viscosity.
Appendix B The relevance of accurate grad-h terms
Figure 7: Effect of the accuracy in grad-h terms. The top row shows the value of the grad-h correction $\Omega$ for all particles in a $250^{3}$ simulation at $t=10$. Left, $\Omega$ is calculated using the density field directly. Right, using the differences in the density field. The red lines show the mean value of $\Omega$. The bottom row shows a thin slice of the resulting density field.
Grad-h terms are an important addition to the SPH technique when using variable smoothing lengths ($h$) (Springel and Hernquist 2002; Monaghan 2002). It takes into account the spatial variation of $h$ as a corrective term in the momentum (Eq. 22) and energy (Eq. 23) equations:
$\Omega_{a}=\left[1+\frac{h_{a}}{3\rho_{a}}\frac{\partial\rho_{a}}{\partial h_{a}}\right]^{-1}\,.$ (24)
These corrections are a standard method to improve accuracy and consistency in particle interpolation, reducing numerical instability in variable-density simulations, and as such, they are included by default in most astrophysics SPH codes.
During the subsonic simulations presented above, we noticed that the density field appeared ”granulated” only when grad-h corrections were used jointly with SLR dissipation (see Fig. 7, left column). Initially, we suspected that the new addition (the SLR dissipation control) was the source of the problem, but it turned out that the problematic term was how the grad-h correction was calculated. After inspecting the values of this correction we noticed that the average was systematically below 1 and with a relatively large spread. For example for the $250^{3}$ simulation, $\left<\Omega\right>=0.98\pm 0.16$. This doesn’t seem much, but it is exactly the kind of ”small” inconsistency that becomes relevant in subsonic turbulence, in particular when the dissipation formulation is designed to stop damping that noise.
In low-Mach turbulence, physical density fluctuations are weak, so any residual inconsistency in the pressure–density–$h$ coupling becomes visible. We can write the grad-h field as:
$\Omega_{a}=\bar{\Omega}+\delta\Omega_{a}\,.$ (25)
Then, expanding the multiplicative term in the SPH equations:
$\frac{1}{\Omega_{a}}=\frac{1}{\bar{\Omega}}\left(1-\frac{\delta\Omega_{a}}{\bar{\Omega}}+\mathcal{O}(\delta\Omega_{a}^{2})\right)$ (26)
The $1/\Omega_{a}$ multiplicative factor that scales the pressure forces has two error sources: a biased mean and a noisy scatter. The first is just a constant scaling affecting the effective compressibility and shifting the global balance slightly. It shows as a systematic drift: slightly different RMS Mach, different effective forcing response, and slightly different spectra cutoff location, but doesn’t introduce particle-scale structure by itself. The second is the actual source of the granularity, since it is a multiplicative noise field coupled to the pressure acceleration. Since $\delta\Omega_{a}$ varies on the inter-particle scale and correlates with local disorder, it acts like a high-$k$ forcing of the compressive jitter:
$\mathbf{a}_{a}^{P}\approx\frac{1}{\bar{\Omega}}\mathbf{a}^{P}_{\Omega=1}-\frac{1}{\bar{\Omega}^{2}}\delta\Omega_{a}\mathbf{a}^{P}_{\Omega=1}+\dots$ (27)
If these corrections are in fact a noisy field, we get a noisy acceleration and particle jitter. This leads to noisy density estimates, noisy smoothing length, and back to noisy grad-h corrections, creating a feedback loop that distorts the field completely. This is usually masked by baseline AV, but becomes apparent when SLR suppresses dissipation in smooth regions.
According to Eq. 24 we need to evaluate $\partial\rho_{a}/\partial h_{a}$ and, as far as we know, the literature always implements it as a direct derivative of the density field. This is well known to have considerable E0 errors, and a typical trick is to include a scalar in the derivative to calculate it using the differences of the field, rather than the field directly. This has been used extensively to yield built-in conservation in SPH and zero derivatives when the field is constant. For example:
$\frac{\partial\rho}{\partial h}=\frac{1}{\phi}\left[\frac{\partial(\phi\rho)}{\partial h}-\rho\frac{\partial\phi}{\partial h}\right]\,$, (28)
where $\phi$ could be any other field or scalar. If we simply choose $\phi=1$:
$\frac{\partial\rho_{a}}{\partial h_{a}}=\sum_{b}V_{b}\rho_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}-\rho_{a}\sum_{b}V_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}\,.$ (29)
The first term on the RHS of Eq. 29 is the common grad-h term that is used to calculate $\Omega_{a}$. The second term in the RHS of Eq. 29 is the additional correction to evaluate grad-h as differences and reduce its noise.
The practical implementation of this term depends on which volume elements are used. In SPH-EXA, we corrected the calculation of $\partial\rho_{a}/\partial h_{a}$, amending formula A17 from García-Senz et al. (2022). Defining:
$\Psi_{a}\equiv\rho_{a}\sum_{b}V_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}\,$, (30)
we obtain:
$\begin{split}\frac{\partial\rho_{a}}{\partial h_{a}}=\left[\frac{\rho_{a}}{\rho^{0}_{a}}-X_{a}W_{aa}(h_{a})\right]&\sum_{b}m_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}-\rho^{0}_{a}\Psi_{a}+\\ \frac{m_{a}}{X_{a}}&\sum_{b}X_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}-\rho_{a}\Psi_{a}\,.\end{split}$ (31)
Note that if this formulation is used, self-contributions of particle $a$ must not be included. There are different ways to implement this correction, as well as different possibilities for the field $\phi$. We opted for the easiest and least invasive methodology.
As a result, Fig. 7 (right column) shows that the average value of $\Omega$ improved considerably, becoming $\left<\Omega\right>=1.00\pm 0.04$, and the density field was recovered.
Appendix C Second-order velocity structure function
Figure 8: Mass-weighted turbulent Mach number as a function of spatial scale for a turbulent box simulation with a resolution of $2000^{3}$ particles. The dotted line indicates the RMS velocity of the simulation. The dashed line is a power-law fit to the turbulent cascade region.
The SPH method is particularly advantageous for a wide range of mass-weighted analysis methods. Most importantly, it removes the need for special tracer particles as the SPH particles themselves serve as tracers. This allows for an easy computation of several quantities like the finite-time Lyapunov exponent and mass-weighted structure functions. The structure function is defined as
${SF}_{2}(d/L)=\left\langle|\@vec{v}(\@vec{r})-\@vec{v}(\@vec{r}+\@vec{d}/L)|^{2}\right\rangle_{\@vec{r}}$ (32)
where d=|d(r)|d=|\@vec{d}(\@vec{r})| is the distance between the two particles in a pair, and we compute the average over many SPH particle pairs, which makes this a mass-weighted structure function. $L$ is the turbulent driving scale. From this, one can compute the turbulent Mach number as
$M_{t}(d/L)=\sqrt{{SF}_{2}(d/L)/(2c_{s}^{2})}$ (33)
As an example, Figure 8 shows the turbulent Mach number as a function of the spatial scale for one of our SPH-EXA simulations. The computation was done using $10^{9}$ particle pairs per snapshot and was averaged over $41$ snapshots of the simulation, which is sufficient to achieve statistical convergence.
In the regime of the turbulent cascade, we find a power law with an exponent of $0.333$. This perfectly matches the expected result of $1/3$ for subsonic Kolmogorov turbulence. Together with the density PDF shown in Figure 4, this demonstrates that SPH-EXA is capable of producing accurate, mass-weighted statistics for subsonic turbulence in a much more natural way than comparable mesh codes.
Appendix D High-resolution view of the velocity field
Figures 9 to 11 show a detailed image of a slice of the velocity magnitude, density, and enstrophy for the $3200^{3}$ simulation, respectively, where we can appreciate the very small-scale features resolved with SPH-EXA. All slices have a thickness of $2h$ and they are centered on $z=0$. Particles are displayed directly (no particle-to-grid interpolation is applied) and each slice contains $\sim 5.84\times 10^{7}$ particles.
Figure 9: Velocity magnitude $|\mathbf{v}|$ at $t=10$ for the $3200^{3}$ simulation, shown in a slice of thickness $2h$ centered on $z=0$.
Figure 10: Density at $t=10$ for the $3200^{3}$ simulation, shown in a slice of thickness $2h$ centered on $z=0$.
Figure 11: Enstrophy at $t=10$ for the $3200^{3}$ simulation, shown in a slice of thickness $2h$ centered on $z=0$.